OpenStudy (anonymous):

homework check... square root of p^7 is p^7/2 right

OpenStudy (anonymous):



OpenStudy (anonymous):

equals p^7/2

zepdrix (zepdrix):

Yes :)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

In general: \[\large \sqrt{b^a} \implies \color{green}{b^{\frac{a}{2}}}\]

OpenStudy (anonymous):

i mean "yes" but that is not what you are supposed to write as a final answer

OpenStudy (anonymous):

\[\sqrt{p^7}=\sqrt{p^6\times p}=\sqrt{p^6}\times \sqrt{p}=p^3\sqrt{p}\]

OpenStudy (anonymous):

ohh well what should i write did i mess up?

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

so satellite73 is correct and mine was not simplified?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

You can do that what you did if you have nothing to solve or factor out or if you have nothing of which you can take square root..

OpenStudy (anonymous):

oh i see i'll try a different homework problem